A small publishing company is planning to publish a new book. The production costs will include one-time fixed costs (such as editing) and variable costs (such as printing). The one-time fixed costs will total $50,560. The variable costs will be $8.50 per book. The publisher will sell the finished product to bookstores at a price of $24.50 per book. How many books must the publisher produce and sell so that the production costs will equal the money from sales?
Found 2 solutions by Zucchini, Theo:
Answer by Zucchini(70)

First, write the equation for the production of the books:
p = 8.50b + 50,560
This equation shows that for the production, $8.50 is paid per book and $50,560 have to paid anyway.
Now, write the equation for the sales:
s = 24.50b
Now, set both of the equations equal to each other because the sales and production money have to be equal, as asked in the question.
p = s
8.50b + 50,560 = 24.50b
50,560 = 16b
3,160 = b
3,160 books must be produced and sold.

Answer by Theo(13342)

x = number of books.

cost = 50,560 + 8.50 * x

revenue = 24.5 * x

break even is when revenue equals cost.

revenue equals cost equation becomes:

24.5 * x = 50,560 + 8.50 * x

subtract 8.50 * x from both sides of the equation to get:

24.5 * x - 8.5 * x = 50,560

combine like terms to get:

16 * x = = 50,560

divide both sides of the equation by 16 to get:

x = 50,560 / 16 = 3160.

replace x in the original cost and revenue equations to get:

cost = 50,560 + 8.50 * x becomes cost = 50,560 + 8.50 * 3160 which gets you cost = 77,420.

revenue = 24.5 * x becomes revenue = 24.5 * 3160 which gets you revenue = 77,420.

solution looks good.

he will have to sell 3160 books to break even.